OpenStudy (anonymous):

So i need help with my chemistry homework. The question is; A sample of copper with a mass of 63.5g contains 60.2 x 10^23 atoms. Calculate the mass of a single copper atom?

OpenStudy (mayankdevnani):

\[\large \bf 63.5 g~copper~contains~60.2 \times 10^{23}~atoms\] so, \[\large \bf 1~atom~weigh~\frac{63.5}{60.2 \times 10^{23}}g=1.054 \times 10^{-23}g\]

OpenStudy (mayankdevnani):

understood? @karinaneedshelp15621

OpenStudy (anonymous):

yes, thanks for the help @mayankdevnani

OpenStudy (mayankdevnani):

welcome :)